solve using the elimination method.
x-y=3
-3x+y=7

Answer by rfer(16322)

x=y+3
-3(y+3)+y=7
-3y-9+y=7
-2y=16
y=-8
-------------
x-y=3
x+8=3
x=-5
---------------
-3x+y=7
-3(-5)-8=7
15-8=7